OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8

OpenStudy (amorfide):

if you rearrange the first equation -x + 2y = -6 to get 2y+6=x you know what x is so you replace all x in the second equation with 2y+6 to get 3(2y+6) + y = 8

OpenStudy (anonymous):

1) -2y + 6 2) 2y - 6 3) 2y + 6 4) -2y - 6

OpenStudy (amorfide):

I just told you

OpenStudy (anonymous):

IK THAT

OpenStudy (anonymous):

thank you can u help me with another one ?